A small producer of machine tools wants to move to a larger building, and has identified two alternatives. Location A has annual fixed costs of $160,000 and variable costs of $15,000 per unit; location B has annual fixed costs of $360,000 and variable costs of $14,000 per unit. The finished items sell for $16,000 each.
Question: what volume of output would the two locations have the same total cost?

Answer by math_tutor2020(3835)


x = number of units produced
y = total cost in thousands of dollars

eg: y = 5 means the total cost is 5000 dollars

Cost equations:
Location A: y = 15x + 160
Location B: y = 14x + 360
The general template is
total cost = (variable cost)*x + (fixed cost)

Since we want the two costs to be the same, we're dealing with the same y value.
Apply substitution and solve for x.

y = 14x+360
15x+160 = 14x+360
15x-14x = 360-160
x = 200
You need to produce 200 units to get the same cost for each location.

Location A:
y = 15x+160
y = 15*200+160
y = 3,000+160
y = 3,160
Location B:
y = 14x+360
y = 14*200+360
y = 2,800+360
y = 3,160
Getting the same result for y helps confirm the correct x value.

Remember that y is the total cost in thousands. So we multiply that y value by 1000 to get 1000*3160 = 3,160,000

Each location will have a total cost of $3,160,000

--------------------------------------------

Answer: 200 units